# CTD4 - A Deep Continuous Distributional Actor-Critic Agent with a Kalman Fusion of Multiple Critics

### v1

@@ 1 Introduction
(Lyle et al., 2019; Rowland et al., 2018; Bellemare et al., 2017; Choi et al., 2019) are some examples of related literature using distributions in RL in discrete control problems, where the output of the Q-function approximator is a categorical distribution. This type of distribution is usually highly dependent on several hyperparameters, such as the number of categories or bounding values. Furthermore, to get acceptable performance and reduce the approximation error, task-specific knowledge with additional steps such as projections, truncations, or auxiliary predictions are needed (Bellemare et al., 2017). Consequently, even when the results are satisfactory against traditional deterministic RL approaches, the computations and the training process on categorical distributions are typically elaborate.
@@
Likewise, either in traditional RL or distributional RL, the overestimation bias is one of the primary obstacles to accurate policy learning. When estimating a Q-function, an overestimation is inevitable due to the imperfect nature of the estimator. That is, the approximation may be significantly higher than the actual value. Overestimation may affect the policy training, making it unstable and sub-optimal (Lan et al., 2020; Fujimoto et al., 2018; Kuznetsov et al., 2020). Prior studies addressed this problem through an ensemble of various Q-approximators in multiple ways. For example, (Fujimoto et al., 2018) takes the minimum over two separate network approximations for continuous control problems. While (Lan et al., 2020) mitigates overestimation by taking the minimum value among several approximators for discrete action spaces. On the other hand, (Kuznetsov et al., 2020) combines multiple categorical distributional approximations, averaging and truncating them to create a single $Q$ estimation. We argue that taking the minimum value among the $Q$ approximations is not the best option since the opposite effect of overestimation could happen, i.e. underestimation, ending up in discouragement of exploration (Lan et al., 2020). Additionally, using the minimum value among the approximations makes the ensemble lose its significance since the other values are ignored. Taking the average among the approximations is not optimal either, especially when working with distributions of returns. Weighting all of the distributions the same when they are not equally good could result in sub-optimal or incorrect approximations.
@@
In this work, to mitigate the issues of categorical distribution and overestimation bias, we present a novel Continuous Distributional Reinforcement Learning algorithm called Continuous Twin Delayed Distributed Deep Deterministic Policy Gradient CTD4. We extend and transform the work of (Fujimoto et al., 2018) TD3, from a standard deterministic actor-critic method to a continuous distributional actor-critic approach, including the simplicity in the implementation with the benefits of ensemble distributions for learning an RL control policy for continuous action problems. We propose using a normal distribution to parameterize the approximation of the return distribution. With the normal distribution described by its mean $\mu$ and standard deviation $\sigma$, the disjoint support issues of categorical distributions and the high dependency on task-specific hyperparameters can be avoided. In the same way, no customized loss/distance metrics are needed. Furthermore, to alleviate the overestimation, we propose an ensemble of the distributional Q-approximators fused through a Kalman method, preserving the power of the ensemble. The paper contributions are:
@@ 4 CTD4
To minimize the TD error, characterized as the discrepancy between distributions in distributional RL, and facilitate the training of the approximators $Z_{\theta n}$ for $n\in[1..N]$, it is imperative to formulate the Bellman target function as a normal distribution. Specifically, we express the target distribution as $Z_{\text{target}}=R+\gamma\ Z(\mu_{k},\sigma_{k})$. Since we are assuming a Gaussian normal distribution $X\sim\mathcal{N}(\mu,\sigma)$, the transformation and ship rule $Y=aX+b$ can be applied; therefore the target distribution can be defined as:
@@ 4.1 Optimal Fusion Strategy: Why a Kalman Fusion?
Utilizing an ensemble of approximations has been a recurring strategy in deterministic and categorical RL methods to address overestimation issues inherent in $Q$ approximations. Common practices involve selecting the minimum value among Q-Values (Fujimoto et al., 2018; Lan et al., 2020) or averaging all the approximations (Kuznetsov et al., 2020) as methods to process the output of the ensemble critics. However, we argue that both of these approaches have limitations. Opting for the minimum value prioritizes a singular value, underutilizing the potential power of the ensemble. On the other hand, taking the average assigns equal significance to all approximations, irrespective of their potential sub-optimality or inaccuracy.
(a)                                    (b)
Figure 4: a) A comparative analysis conducted among minimum value, average, and Kalman fusion methods. b) A comparative analysis conducted in order to determine the optimal ensemble size. The results are averaged across five independent seeds, with the shaded area representing the standard deviation.
Notably, in our proposed framework, the output of each critic is a normal distribution parameterized by $\mu$ and $\sigma$. This perspective allows us to interpret each approximation as a sensor reading, enabling the application of sensor fusion methods. To enhance Q-value estimation and fully leverage the collective strength of the ensemble, we employ a Kalman filter to fuse multiple distribution approximations. The Kalman fusion identifies the most probable approximation within the ensemble by calculating the joint probability of the distributions. This strategic fusion optimises Q-value estimates and endows the RL agent with resilience to noise, errors, or inaccuracies inherent in the individual approximations.
It is crucial to emphasize that we do not keep track of all the previous approximations; we only fuse them in each training step. In Figure 4 a), a comparative analysis is presented. Under identical conditions, we conducted an experiment, altering only the ensemble fusion method among Kalman, minimum value, and average fusion. The results indicate that our proposed method, employing Kalman fusion, exhibits greater stability and achieves higher rewards with fewer samples compared to other fusion methods.
@@ 4.2 Ensemble Size Optimization: Determining the Optimal Number of Critics
We conducted an analysis to comprehend the influence of ensemble size. Increasing the number of critics can significantly mitigate overestimation, leading to a potential enhancement in overall performance. In other words, a higher number of approximations corresponds to a more pronounced alleviation of overestimation (Lan et al., 2020; Kuznetsov et al., 2020). Nevertheless, if the number of critics becomes excessively high, there is an elevated risk of instability, potentially leading to underestimation, discouraging exploration, and, consequently, yielding a suboptimal policy (Lan et al., 2020). Additionally, the computational cost experiences a substantial increase with the growing number of critics in the ensemble. To decide the correct ensemble size, we train the agent under the same conditions, fusing the approximation with a Kalman fusion but varying only the number of approximators. Figure 4 b) shows the analysis result where it is easy to confirm the idea proposed above. A suboptimal performance may occur if the number of critics is too high (for instance, N= 10). An ensemble of three was selected based on these results for the rest of the experiments.
@@ 5 Experiments
Our experimental testing aims to understand the effects and efficiency of Continuous Distributional approximators for continuous action spaces. We conduct comparative analyses against TD3, utilizing the official source-code implementation provided by the author  [^1] https://github.com/sfujim/TD3.
We purposefully avoid applying or comparing our method with previous approaches, such as C51 or IQN, which are tailored for discrete action spaces. Our research squarely focuses on addressing the challenges posed by complex tasks characterized by continuous action spaces—a domain notorious for its heightened difficulty level. Furthermore, given TD3’s established dominance over a spectrum of algorithms designed for continuous action spaces, our comparative analysis is exclusively directed towards evaluating our method against TD3. To conduct experiments with our proposed method, we select the challenging continuous control tasks from the DeepMind Control Suite (DMCS) (Tassa et al., 2018). DMCS is an established and widely recognized RL test-bed, offering complex environments custom-made for continuous control tasks. We choose ten environments, each presenting distinct levels of complexity and task requirements. See Figure 6 for a graphic representation of each environment.
@@ 6 Results and Discussion
Figure 5: Learning curves for CTD4 and TD3 on ten DeepMind Control Suite continuous control tasks. Our proposal is superior in nine of the ten tasks evaluated against TD3. In each plot, the solid line represents the average, while the shaded region represents the standard deviation of the average evaluation over five independent seeds. Curves have not been smoothed to demonstrate clarity in our results
@@
Following thorough testing and analysis, the results presented in Figure 5 illustrate the average evaluation reward across multiple seeds. Each plot serves to visually convey and compare the performance of our proposed CTD4 against the baseline TD3 algorithm under identical conditions.
@@
Using a continuous distribution approach parametrized by the mean and standard deviation, coupled with the fusion of multiple critics through a Kalman method, is effective in achieving comparable or enhanced performance across diverse tasks. The superiority of our proposed methodology becomes evident, particularly in complex scenarios where TD3 struggled to solve tasks or exhibit patterns of task completion. Our proposal demonstrates superior performance, outperforming TD3 in nine of the ten tests conducted.
@@
It is crucial to note that we deliberately selected exceptionally challenging tasks to assess the performance of our proposal. Specific tasks, such as Reacher Hard, Finger Spin Hard, or Ball-in-Cup, pose increased difficulty due to their sparse reward structures, adding complexity to task-solving. Conversely, tasks like Walker Walk and Humanoid Run demand intricate coordination among numerous joints, making them inherently complex. The Acrobot Swingup task necessitates balance and precise control, while the Fish Swim task involves navigating a 3D environment, introducing an additional layer of complexity to the learning process. The results affirm that the stochastic nature of CTD4 can discover superior solutions with higher rewards. This becomes particularly evident in environments such as Fish Swim or Acrobot Swingup, where the performance curves demonstrate that our proposal can outperform TD3 without the need for complex hyperparameter tuning or intricate projection steps as required by other distributed RL methods.
@@
However, in the spirit of transparency regarding the results, it is important to acknowledge that in a specific scenario, Chetah Run, our proposal employing continuous distributions fails to surpass TD3. While the performance is reasonably similar, TD3 exhibits a marginal superiority in this particular case. We believe that this may arise due to distinctive behaviours exhibited by Cheetah Run compared to other environments in the test bed. There could be subtleties in the reward distribution relative to certain actions that were not explicitly considered in our approach. Further investigation is warranted to comprehend the specific dynamics at play in Cheetah Run, which might require tailored adjustments to enhance the performance of our proposal in this particular setting.
@@ Appendix A Hyperparameters
Table 1: Complete list of Hyperparameters employed in our proposal. TD3 values are based on the official implementation provided by the author.
TD3(Fujimoto et al., 2018) Ours
Hyperparameter Value Value
Actor Learning Rate $3.10^{-4}$ $10^{-4}$
Critic Learning Rate $3.10^{-4}$ $10^{-3}$
@@
Decrease Noise per Step —-

### v2

@@ 1 Introduction
(Lyle et al., 2019; Rowland et al., 2018; Bellemare et al., 2017; Choi et al., 2019) are some examples of related literature using distributions in RL in discrete control problems, where the output of the Q-function approximator is a categorical distribution. This type of distribution is usually highly dependent on several hyperparameters, such as the number of categories or bounding values. Furthermore, to get acceptable performance and reduce the approximation error, task-specific knowledge with additional steps such as projections, truncations, or auxiliary predictions are needed (Bellemare et al., 2017). Consequently, even when the results are satisfactory against traditional deterministic RL approaches, the computations and the training process on categorical distributions are typically complicated.
@@
Likewise, either in traditional RL or distributional RL, the overestimation bias is one of the primary obstacles to accurate policy learning. When estimating a Q-function, an overestimation is inevitable due to the imperfect nature of the estimator. That is, the approximation may be significantly higher than the actual value. Overestimation may affect the policy training, making it unstable and sub-optimal (Lan et al., 2020; Fujimoto et al., 2018; Kuznetsov et al., 2020). Prior studies addressed this problem through an ensemble of various Q-approximators in multiple ways. For example, (Fujimoto et al., 2018) takes the minimum over two separate network approximations for continuous control problems. While (Lan et al., 2020) mitigates overestimation by taking the minimum value among several approximators for discrete action spaces. (Kuznetsov et al., 2020) combines multiple categorical distributional approximations, averaging and truncating them to create a single $Q$ estimation. While (Chen et al., 2021) combines ensemble learning and double Q-learning principles to improve stability by averaging predictions from the ensemble.
We argue that taking the minimum value among the $Q$ approximations is not the best option since the opposite effect of overestimation could happen, i.e. underestimation, ending up in discouragement of exploration (Lan et al., 2020). Additionally, using the minimum value among the approximations makes the ensemble lose its significance since the other values are ignored. Taking the average among the approximations is not optimal either, especially when working with distributions of returns. Weighting all of the distributions the same when they are not equally good could result in sub-optimal or incorrect approximations.
@@
In this work, to mitigate the issues of categorical distribution and overestimation bias, we present a novel Continuous Distributional Reinforcement Learning algorithm called Continuous Twin Delayed Distributed Deep Deterministic Policy Gradient CTD4. We extend and transform the work of (Fujimoto et al., 2018) TD3, from a standard deterministic actor-critic method to a continuous distributional actor-critic approach, including the simplicity in the implementation with the benefits of ensemble distributions for learning an RL control policy for continuous action problems. We propose using a normal distribution to parameterize the approximation of the return distribution. With the normal distribution described by its mean $\mu$ and standard deviation $\sigma$, the disjoint support issues of categorical distributions and the high dependency on task-specific hyperparameters can be avoided. In the same way, no customized loss/distance metrics are needed. Furthermore, to alleviate the overestimation, we propose an ensemble of the distributional Q-approximators fused through a Kalman method, preserving the power of the ensemble without the need for complex tuning parameters. The paper contributions are:
@@ 4 CTD4
To minimize the TD error, characterized as the discrepancy between distributions in distributional RL, and facilitate the training of the approximators $Z_{\theta n}$ for $n\in[1..N]$, it is imperative to formulate the Bellman target function as a normal distribution. Specifically, we express the target distribution as $Z_{\text{target}}=R+\gamma\ Z(\mu_{k},\sigma_{k})$. Since we are assuming a Gaussian normal distribution $X\sim\mathcal{N}(\mu,\sigma)$, the linear Gaussian transformation rule $Y=aX+b$ can be applied; therefore the target distribution can be defined as:
@@ 4.1 Optimal Fusion Strategy: Why a Kalman Fusion?
Utilizing an ensemble of approximations has been a recurring strategy in deterministic and categorical RL methods to address overestimation issues inherent in $Q$ approximations. Common practices involve selecting the minimum value among Q-Values (Fujimoto et al., 2018; Lan et al., 2020) or averaging all the approximations (Kuznetsov et al., 2020; Chen et al., 2021) as methods to process the output of the ensemble critics. A clear example of this is REDQ which averages the predictions from the ensemble. However, REDQ is limited to deterministic setups and cannot handle distributional updates, whether categorical or continuous. Moreover, its high update rate (G=20) significantly escalates computational costs and risks overfitting without diverse training data. Additionally, although employing an ensemble approach, REDQ only employs two critics randomly selected from the ensemble, potentially underutilizing the ensemble’s power.
We argue that these approaches have limitations. Opting for the minimum value prioritizes a singular value, underutilizing the potential power of the ensemble. On the other hand, taking the average assigns equal significance to all approximations, irrespective of their potential sub-optimality or inaccuracy.
Figure 4: Fusion Method Analysis: Kalman Fusion (Proposed) versus Minimum and Average Value Methods in two environments.
Notably, in our proposed framework, the output of each critic is a normal distribution parameterized by $\mu$ and $\sigma$. This perspective allows us to interpret each approximation as a sensor reading, enabling the application of sensor fusion methods. To enhance Q-value estimation and fully leverage the collective strength of the ensemble, we employ a Kalman filter to fuse multiple distribution approximations. The Kalman fusion identifies the most probable approximation within the ensemble by calculating the joint probability of the distributions. This strategic fusion optimises Q-value estimates and endows the RL agent with resilience to noise, errors, or inaccuracies inherent in the individual approximations. It is crucial to emphasize that we do not keep track of all the previous approximations; we only fuse them in each training step.
Under identical conditions, only the ensemble fusion method is altered among Kalman fusion, minimum value fusion, and average fusion. The agent is then trained for a fixed number of steps. The results for two complex tasks from the DMCS (Cheetah run and Cartpole Swingup) are displayed in Figure 4. Each fusion method for each task was trained using five different seeds. The results indicate that the proposed method, employing Kalman fusion, exhibits greater stability and achieves higher rewards with fewer samples compared to other fusion methods.
@@ 4.2 Ensemble Size Optimization: Determining the Optimal Number of Critics
Figure 5: Ensemble Size Analysis Varying Number of Critics Under Identical Conditions.
@@
We conducted an analysis to comprehend the influence of ensemble size. Increasing the number of critics can significantly mitigate overestimation, leading to a potential enhancement in overall performance. In other words, a higher number of approximations corresponds to a more pronounced alleviation of overestimation (Lan et al., 2020; Kuznetsov et al., 2020). Nevertheless, if the number of critics becomes excessively high, there is an elevated risk of instability, potentially leading to underestimation, discouraging exploration, and, consequently, yielding a suboptimal policy (Lan et al., 2020). Additionally, the computational cost experiences a substantial increase with the growing number of critics in the ensemble. To determine the correct ensemble size, an agent is trained under the same conditions, using a Kalman fusion for the approximation but varying only the number of approximators. The analysis results, conducted across five different seeds for two complex unrelated tasks (Reacher Hard and Cartpole Swingup), are illustrated in Figure 5. Suboptimal performance may occur if the number of critics is too high (for instance, $N=10$). Based on these results, an ensemble $N=3$ was selected.
@@ 5 Experiments
Our experimental testing aims to understand the effects and efficiency of Continuous Distributional approximators for continuous action spaces. We conduct comparative analyses against TD3 and REDQ, utilizing the official source-code implementation provided by the authors. REDQ is selected for its similarity to using an ensemble method, aligning with the approach taken in this chapter. TD3 is chosen for its well-established dominance across a range of algorithms designed for continuous action spaces.
We purposefully avoid applying or comparing our method with previous approaches, such as C51 or IQN, which are tailored for discrete action spaces. Our research squarely focuses on addressing the challenges posed by complex tasks characterized by continuous action spaces—a domain notorious for its heightened difficulty level.
To conduct experiments with our proposed method, we select the challenging continuous control tasks from the DeepMind Control Suite (DMCS) (Tassa et al., 2018). DMCS is an established and widely recognized RL test-bed, offering complex environments custom-made for continuous control tasks. We choose ten environments, each presenting distinct levels of complexity and task requirements. See Figure 7 for a graphic representation of each environment.
@@ 6 Results and Discussion
Figure 6: Learning curves for CTD4, REDQ and TD3 on ten DeepMind Control Suite continuous control tasks. In each plot, the solid line represents the average, while the shaded region represents the standard deviation of the average evaluation over five independent seeds. Curves have not been smoothed to demonstrate clarity in our results
@@
Following thorough testing and analysis, the results presented in Figure 6 illustrate the average evaluation reward across multiple seeds. Each plot serves to visually convey and compare the performance of our proposed CTD4 against TD3 and REDQ under identical conditions.
@@
Using a continuous distribution approach parametrized by the mean and standard deviation, coupled with the fusion of multiple critics through a Kalman method, is effective in achieving comparable or enhanced performance across diverse tasks. The superiority of our proposed methodology becomes evident, particularly in complex scenarios where TD3 and REDQ struggled to solve tasks or exhibit patterns of task completion.
@@
It is crucial to note that we deliberately selected exceptionally challenging tasks to assess the performance of our proposal. Specific tasks, such as Reacher Hard, Finger Spin Hard, or Ball-in-Cup, pose increased difficulty due to their sparse reward structures, adding complexity to task-solving. Conversely, tasks like Walker Walk and Humanoid Run demand intricate coordination among numerous joints, making them inherently complex. The Acrobot Swingup task necessitates balance and precise control, while the Fish Swim task involves navigating a 3D environment, introducing an additional layer of complexity to the learning process. The results affirm that the stochastic nature of CTD4 can discover superior solutions with higher rewards. This becomes particularly evident in environments such as Hopper Hop or Acrobot Swingup, where the performance curves demonstrate that our proposal can outperform TD3 and REDQ without the need for complex hyperparameter tuning or intricate projection steps as required by other distributed RL methods.
However, in the spirit of transparency regarding the results, it is important to acknowledge that in a specific environment, “Cheetah Run,” the proposed method employing continuous distributions fails to surpass TD3. While the performance is reasonably similar, TD3 exhibits a marginal superiority in this particular case. A similar scenario occurs with the “Finger Turn Hard” environment, where REDQ narrowly surpasses CTD4. However, TD3 shows the worst performance in this setting.
@@
This discrepancy may arise due to distinctive behaviours exhibited by Cheetah Run and Finger Turn Hard compared to other environments in the test bed. Subtleties in the reward distribution relative to certain actions might not have been explicitly considered in the approach. Further investigation is warranted to understand the specific dynamics at play in these two environments, which might require tailored adjustments to enhance the performance of the proposed method in these particular settings.
@@ Appendix A Hyperparameters
Table 1: Complete list of Hyperparameters employed in our proposal. TD3 and REDQ values are based on the official implementation provided by the author.
TD3(Fujimoto et al., 2018) Ours REDQ(Chen et al., 2021)
Hyperparameter Value Value Value
Actor Learning Rate $3.10^{-4}$ $10^{-4}$ $3.10^{-4}$
Critic Learning Rate $3.10^{-4}$ $10^{-3}$ $3.10^{-4}$
@@
Decrease Noise per Step